OpenStudy (anonymous):

I need to find radius of convergence of: \[\sum_{n=1}^{\infty}\frac{ (-1)^{n+1} }{ 2n }x^n\] Can this be done by? \[\lim_{n \rightarrow \infty}\left| \frac{ a_{n+1} }{ a_n } \right|\]

OpenStudy (anonymous):

Where \[a_n=\frac{ (-1)^{n+1} }{ 2n }\] right?

OpenStudy (anonymous):

It is bugging for me, dont know if it is for you. But to easier see the question. I need to find radius of convergence of: \[\sum_{n=1}^{\infty}\frac{ (-1)^{n+1} }{ 2n }x^n\] Can this be done by? \[\lim_{n \rightarrow \infty}\left| \frac{ a_{n+1} }{ a_n } \right|\]

OpenStudy (anonymous):

yes, use the ratio test

OpenStudy (anonymous):

you get that the limit is 1 pretty much in your head

OpenStudy (anonymous):

and since you are working with the absolute value, you can safely ignore the \((-1)^n\) part

OpenStudy (anonymous):

Is the ratio test \[\lim_{n \rightarrow \infty}\left| \frac{ a_n }{ a_{n+1} } \right|\] or the 1 i posted before?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

Thats what I thought, but then I found the other formular here. https://en.wikipedia.org/wiki/Radius_of_convergence#Theoretical_radius

OpenStudy (anonymous):

but it is only \[\lim_{n\to \infty}\frac{2n+2}{2n}\]

OpenStudy (anonymous):

oh sorry, you are right i had it upside down

OpenStudy (anonymous):

Yeap, got it. Thanks for the help :)